OpenStudy (anonymous):

Consider the given curves to do the following x=4sqrt(y) x=0 y=1 Use the method of cylindrical shells to find the volume V of the solid obtained by rotating the region bounded by the given curves about the x-axis.

OpenStudy (dumbcow):

A(y) = 2pi*y*(4sqrt(y)) \[V=2\pi \int\limits_{0}^{1}4y \sqrt{y} dy\] \[=8\pi \int\limits_{0}^{1} y ^{3/2} dy\] \[=\frac{16}{5}\pi\]